Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $108.00 for 3 days and 300 miles, while Mary was charged $193.00 for 5 days and 600 miles. What does Best Rentals charge per day and per mile?

Answer by stanbon(75887)

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $108.00 for 3 days and 300 miles, while Mary was charged $193.00 for 5 days and 600 miles. What does Best Rentals charge per day and per mile?
------------------------
Equations:
3d + 300m = 108
5d + 600m = 193
--------------------------
Modify for elimination:
6d + 600m = 216
5d + 600m = 193
-----------------------
Subtract and solve for "d"::
d = $23 (daily fee)
-----
Solve for "m":
3d + 300m = 108
69 + 300m = 108
300m = $39
m = 13 cents per mile
============
Cheers,
Stan H.
-------------
==============